Use the square root property to solve the equation:
5z^2 + 5 = 725
Please show me the steps for this. Thank you.

Answer by sachi(548)

5z^2 + 5 = 725
or 5z^2 =725-5=720
or z^2=720/7=144=12^2
or z=12